iEnvyKx:

Say EG = 2 and EF = 8, find the length of segment ED.

iEnvyKx:

1 attachment
Florisalreadytaken:

this is all about proportions we are given that \(\overline{EG} = 2 \ \ \ \vee \ \ \ \overline{EF} = 8\) first lets find \( \overline{GF} \ \ \ | \ \ \ \overline{GF} = \overline{EF} - \overline{GF} = 8-2=6 \ units\) these are similar triangles right,so all the sides are proportional, which means we can set a proportion: \( \frac{ \overline{GD}}{ \overline{EG}}=\frac{ \overline{GF}}{ \overline{GD}} \) \( \frac{ \overline{GD}}{2}=\frac{ 6}{ \overline{GD}} \) \( (\overline{GD} )^2 = 12 \implies \ \overline{GD} = \sqrt{12 } \) so we have this right triangle as drawn  Can you find C?